OpenStudy (anonymous):

Factor x^3+8x^2+2X+16

OpenStudy (anonymous):

ok I know this give me one sec

OpenStudy (luigi0210):

*waits a sec*

OpenStudy (luigi0210):

Welcome to Openstudy :)

OpenStudy (anonymous):

(-1+x) (1+x) (1+x^2)

jhonyy9 (jhonyy9):

x^3 +8x^2 +2x +16 = so try grouping it and you see that from the first 2 terms factorizing out x^2 will get x^2(x+8) and factorizing out the 2 from the 3rd and 4rd term will get 2(x+8) rewriting x^3 +8x^2 +2x +16 = x^2(x+8) +2(x+8) = so because (x+8) is common you can factorizing out and will get the right answer x^3 +8x^2 +2x +16 = x^2(x+8) +2(x+8) = (x+8)(x^2 +2) hope so much that this is understandably sure right easy

OpenStudy (goformit100):

"Welcome to OpenStudy. I can guide regarding this useful site; ask your doubts from me, for it you can message me. Please use the chat for off topic questions. And remember to give the helper a medal, by clicking on "Best Answer". We follow a code of conduct, ( http://openstudy.com/code-of-conduct ). Please take a moment to read it."